# Large Language Model for Vulnerability Detection and Repair: Literature Review and Roadmap

### v1

@@ 4. RQ1: What LLMs have been utilized?
In general, out of the 36 included studies, we identified 15 distinct LLMs that have been utilized.
LLMs Used for Vulnerability Detection. Fig. 3 illustrates the distribution of the LLMs utilized for vulnerability detection. CodeBERT (Feng et al., 2020) emerges as the predominant LLM in addressing vulnerability detection to date, representing 38.5% (15/39) of the use of LLMs in the included studies. Following closely, GPT-3.5 becomes the second most frequently studied model, accounting for 10.3% (4/39). Regarding the categories of LLMs, encoder-only LLMs comprise 61.5% (24/39) of the use of LLMs in the included studies. Following that, decoder-only LLMs account for 28.2% (11/39), and encoder-decoder LLMs constitute 10.3% (4/39).
LLMs Used for Vulnerability Repair. Fig. 4 illustrates the distribution of LLMs utilized for vulnerability repair. Unlike the detection task, CodeT5 (Wang et al., 2021) emerges as the predominant LLM in addressing vulnerability repair to date, representing 24.0% (6/25) of the use of LLMs in the included studies. Following closely, domain-specific pre-trained Seq2Seq Transformers become the second most frequently studied model, accounting for 16.0% (4/25) of the use of LLMs. Regarding the categories of LLMs, encoder-decoder LLMs comprise 48% (12/25) of LLMs used. Following that, decoder-only LLMs and encoder-only LLMs constitute 44% (11/25) and 8% (2/25) of the use of LLMs, respectively.
Answer to RQ1: Our analysis indicates that, to date, encoder-only LLMs have dominated vulnerability detection, while encoder-decoder LLMs have been prominent for vulnerability repair.
@@ 5. RQ2: How are LLMs adapted for vulnerability detection?
In RQ2, we shift our focus to examining the specific adaptation techniques of LLMs in vulnerability detection. Regarding the usage of LLMs, we summarize three major categories from the included studies: 1) fine-tuning (Devlin et al., 2018), which updates the parameters of LLMs using a labeled dataset, 2) zero-shot prompting (Brown et al., 2020), which freezes the model parameters and does not consider any labeled data, and 3) few-shot prompting (Brown et al., 2020), which also freezes the parameters but considers a few labeled example data points. As depicted in Fig. 5, 82% of the studies utilize fine-tuning, while 11% and 7% employ zero-shot prompting and few-shot prompting, respectively.
@@
In the following subsection, we will introduce detailed categories of LLM adaptation techniques for each of the following: 1) fine-tuning, 2) zero-shot prompting, and 3) few-shot prompting.
5.1. How are LLMs adapted for Vul. Detection?
5.1.1. Fine-tuning
@@
Data-centric Innovations. Data-centric innovations focus on optimizing the vulnerability detection data used for fine-tuning LLMs. Prior studies (Yang et al., 2023; Croft et al., 2023; Nie et al., 2023; Chen et al., 2023a; Kuang et al., 2023) have found that existing vulnerability detection data can suffer from imbalanced label distribution, noisy or incorrect labels, and scarcity of labeled data. Researchers (Yang et al., 2023; Wen et al., 2023; Kuang et al., 2023) have explored how to address the data issues.
@@
  • Data Sampling: To address label imbalance, i.e., having more non-vulnerable code samples than vulnerable ones in the dataset, Yang et al. (Yang et al., 2023) applied various data sampling techniques. They found that random oversampling on raw code data enhances the ability of the LLM-based vulnerability detection approach to learn real vulnerable patterns.
@@
  • Positive and Unlabeled Learning: To address label quality issues such as noisy or incorrect labels, Wen et al. (Wen et al., 2023) proposed PILOT, which learns solely from positive (vulnerable) and unlabeled data for vulnerability detection. Specifically, PILOT generates pseudo-labels for selected unlabeled data and mitigates the data noise by using a mixed-supervision loss.
@@
  • Counterfactual Training: To enhance the diversity of labeled data, Kuang et al. (Kuang et al., 2023) proposed perturbing user-defined identifiers in the source code while preserving the syntactic and semantic structure. This approach generates diverse counterfactual training data, which refers to hypothetical data (e.g., data after perturbing identifiers) differing from actual data (i.e., data without perturbation), useful for analyzing the effect of certain factors. Incorporating these counterfactual data enriches the training data for LLMs.
@@
Combination of LLM with Program Analysis. Many LLMs undergo pre-training on extensive datasets through unsupervised objectives like masked language modeling (Devlin et al., 2018) or next token prediction (Radford et al., 2019). For those LLMs, they may prioritize capturing sequential features and could potentially overlook certain structural aspects crucial for understanding code. To address this limitation, several studies have proposed integrating program analysis techniques with LLMs. The idea involves utilizing program analysis to extract structural features/relations within code, which are then incorporated into LLMs to enhance their understanding. Specifically, Liu et al. (Liu et al., 2024) leveraged Joern (Yamaguchi, 2023) to build the AST and PDG of the function, leveraging this data to pre-train their LLM to predict statement-level control dependencies and token-level data dependencies within the function. Peng et al. (Peng et al., 2023) utilized program slicing to extract control and data dependency information, aiding LLMs in vulnerability detection. Wang et al. (Wang et al., 2024b) proposed to learn program presentations by feeding static source code information and dynamic program execution traces with LLMs. Additionally, Zhang et al. (Zhang et al., 2023c) proposed decomposing syntax-based Control Flow Graphs (CFGs) into multiple execution paths and feeding these paths to LLMs for vulnerability detection.
@@
Combination of LLM with Other Deep Learning Modules. LLMs have their own inherent limitations. Firstly, most LLMs are based on the Transformer architecture, which primarily models sequential relations and features. Secondly, some LLMs (e.g., CodeBERT) impose restrictions on the length of input code snippets. For instance, the most frequently used LLM in vulnerability detection, CodeBERT, can only process 512 tokens. To address these limitations, researchers (Tang et al., 2023; Ziems and Wu, 2021) have attempted to combine LLMs with other DL modules:
  • LLM+GNN: To leverage the structural features of code more effectively, Tang et al. (Tang et al., 2023) introduced CSGVD, which employs graph neural networks (GNN) to extract the graph features of code and combine them with the features extracted by CodeBERT.
@@
  • LLM+Bi-LSTM: To address the length constraints of LLMs, Ziems et al. (Ziems and Wu, 2021) first segmented the input code into multiple fixed-size segments. They then utilized BERT to encode each segment and incorporated a Bidirectional Long Short-Term Memory (Bi-LSTM) module to process the output of BERT on each segment. Finally, a softmax classifier was applied to the last hidden state of the Bi-LSTM to produce the final classification scores.
@@
Domain-specific Pre-training. Domain-specific pre-training involves pre-training an LLM on data specific to a particular domain, such as vulnerability data, before fine-tuning it for a specific task within that domain. This process enables the LLM to better understand the data relevant to the domain. Several studies in the field of vulnerability detection have adopted this technique (Hanif and Maffeis, 2022; Liu et al., 2024; Ni et al., 2023; Wang et al., 2024b). These studies typically use one of three pre-training objectives:
@@
  • Masked Language Modeling: This pre-training objective involves training the LLM to predict masked tokens in a corrupted code snippet. Hanif and Mahmood (Hanif and Maffeis, 2022) introduced VulBERTa, which pre-trained a RoBERTa (Liu et al., 2019) model on open-source C/C++ projects using the Masked Language Modeling objective. Since C/C++ is the programming language of the vulnerability detection data used to evaluate VulBERTa, this pre-training enhances VulBERTa’s ability to understand C/C++ code.
@@
  • Contrastive Learning: This pre-training objective is to minimize the distance between similar functions while maximizing the distance between dissimilar functions. Through contrastive learning, the LLM can learn to capture the distinctive features of code. In particular, Ni et al. (Ni et al., 2023) and Wang et al. (Wang et al., 2024b) utilized different hidden dropout masks to transform the same input function into positive samples (i.e., those similar to the input function) while regarding other distinct functions as dissimilar samples. Then they pre-trained an LLM following the contrastive learning objective.
@@
  • Predicting Program Dependencies: This pre-training objective aims to enhance the LLM by guiding the model to learn the knowledge necessary for analyzing dependencies in programs. Specifically, Liu et al. (Liu et al., 2024) pre-trained an encoder-only LLM on code from open-source C/C++ projects to predict the statement-level control dependency and token-level data dependency.
@@
Causal Learning. Although promising, LLMs have been found to lack robustness under perturbation or when encountering out-of-distribution (OOD) data (Mahbubur Rahman et al., 2024). Rahman et al. (Mahbubur Rahman et al., 2024) suggested that this weak robustness may be attributed to LLMs learning non-robust features, such as variable names, that have spurious correlations with labels. To address this issue, Rahman et al. proposed CausalVul, which first designs perturbations to identify spurious features and then applies causal learning algorithms, specifically do-calculus, on top of LLMs to promote causal-based prediction, enhancing the robustness of LLMs in vulnerability detection.
Figure 6. Effectiveness of Adaptation Techniques for Vulnerability Detection
5.1.2. Zero-shot Prompting
For zero-shot prompting, we only observe one adaptation technique (i.e., prompt engineering).
Prompt Engineering. Researchers (Fu et al., 2023; Zhou et al., 2024b; Purba et al., 2023) have attempted to devise effective prompts to guide LLMs in conducting vulnerability detection in the zero-shot prompting manner. Specifically, Zhou et al. (Zhou et al., 2024b) devised a prompt: “Now you need to identify whether a method contains a vulnerability or not. If has any potential vulnerability, output: ‘this code is vulnerable’. Otherwise, output: ‘this code is non-vulnerable’. The code is [code]. Let’s start:”. Fu et al. (Fu et al., 2023) proposed the prompt: “Predict Whether the C/C++ function below is vulnerable. Strictly return 1 for a vulnerable function and 0 for a non-vulnerable function without further explanation.” Purba et al. (Purba et al., 2023) devised a prompt: “[code] Is this code vulnerable? Answer in only Yes or No”. In these prompts, “[code]” refers to the input code.
5.1.3. Few-shot Prompting
For few-shot prompting, we only find one adaptation technique (i.e., retrieval augmentation).

### v3

@@ 4. RQ1: What LLMs have been utilized?
In general, out of the 58 included studies, we identified 37 distinct LLMs that have been utilized.
LLMs Used for Vulnerability Detection. Fig. 4 illustrates the distribution of the LLMs utilized for vulnerability detection. Please note that we count the occurrence of LLMs used in each study, and a single study may utilize multiple LLMs. CodeBERT (feng2020codebert, ) emerges as the predominant LLM in addressing vulnerability detection to date, representing 26.1% (24/92) of all the use of LLMs in the included studies. GPT-3.5 becomes the second most frequently studied model, accounting for 9.8% (9/92). Regarding the categories of LLMs, encoder-only LLMs comprise 47.8% (44/92) of the use of LLMs in the included studies. Following that, decoder-only LLMs account for 23.9% (22/92), and encoder-decoder LLMs constitute 9.8% (9/92). Lastly, commercial LLMs with undisclosed architectures, such as GPT-3.5 and GPT-4, account for 18.5% (17/92) of the LLM usages.
LLMs Used for Vulnerability Repair. Fig. 5 illustrates the distribution of LLMs utilized for vulnerability repair. Unlike the detection task, CodeT5 (wang2021codet5, ) emerges as the predominant LLM in addressing vulnerability repair to date, representing 15.6% (7/45) of the use of LLMs in the included studies. GPT-3.5 becomes the second most frequently studied model, accounting for 11.1% (5/45). Following closely, GPT-4 and domain-specific pre-trained Seq2Seq Transformers become the third most frequently studied model, accounting for 8.9% (4/45) of the use of LLMs. Regarding the categories of LLMs, decoder LLMs comprise 31.1% (14/45) of LLMs used. Following that, encoder-decoder LLMs and encoder-only LLMs constitute 26.7% (12/45) and 6.7% (3/45) of the use of LLMs, respectively. Lastly, commercial LLMs with undisclosed architectures, such as GPT-3.5 and GPT-4, account for 35.6% (16/45) of the LLM usages.
Table 1. Top 10 LLMs Used in Vulnerability Detection and Repair
Usages Ranking Vulnerability Detection Vulnerability Repair
LLM Structure Size LLM Structure Size
1 CodeBERT Encoder-only 125M CodeT5 Encoder-Decoder 220M
2 GPT-3.5 Unknown Unknown GPT-3.5 Unknown Unknown
3 GPT-4 Unknown Unknown GPT-4 Unknown Unknown
4 CodeT5 Encoder-Decoder 220M Seq2seq Encoder-Decoder 60M
5 UnixCoder Encoder-only 126M CodeGen Decoder-only
350M, 2.7B,
6.1B, 16.1B
6 GraphCodeBERT Encoder-only 125M Codex Unknown 12.0B
7 CodeLlama Decoder-only
7B, 13B,
34B, 70B
Incoder Decoder-only 1.3B, 6.7B
8 VulBERTa Encoder-only 125M PLBART Encoder-Decoder 406M
9 RoBERTa Encoder-only 125M UnixCoder Encoder-only 126M
10 BERT Encoder-only 109M CodeGPT Decoder-only 124M
In addition, Table 1 presents the architecture and model sizes of the top 10 most commonly used LLMs in vulnerability detection and repair. The architectures and model sizes of GPT-3.5 and GPT-4, along with the architecture of Codex, remain undisclosed due to their commercial status. As shown in Table 1, the LLMs widely utilized for vulnerability detection are predominantly lightweight models, with many having sizes under or equal to 126 million parameters. Most of these models are encoder-only architectures. In contrast, the models commonly used for vulnerability repair tend to be larger, with several exceeding 1 billion parameters, such as CodeGen (DBLP:conf/iclr/NijkampPHTWZSX23, ) and Incoder (DBLP:conf/iclr/FriedAL0WSZYZL23, ). The majority of these models have more than 126 million parameters and are primarily decoder-only and encoder-decoder architectures.
Answer to RQ1: Our analysis indicates that, to date, encoder-only LLMs have dominated vulnerability detection, while commercial LLMs and decoder-only LLMs have been prominent in vulnerability repair.
@@ 5. RQ2: How Are LLMs Adapted for Vulnerability Detection?
In RQ2, we shift our focus to examining the specific adaptation techniques of LLMs in vulnerability detection. Regarding the usage of LLMs, we summarize three major categories from the included studies: 1) fine-tuning (devlin2018bert, ), which updates the parameters of LLMs using a labeled dataset, 2) prompt engineering (brown2020language, ), which designs prompts to guide LLMs in generating relevant responses without updating parameters, and 3) retrieval augmentation (RAG) (DBLP:conf/nips/LewisPPPKGKLYR020, ), which integrates knowledge from retrieval systems into the LLMs’ context to improve their performance, also without changing LLMs’ parameters. As depicted in Fig. 6, 73% of the studies utilize fine-tuning, while 17% and 10% employ prompt engineering and retrieval augmentation, respectively.
@@
Table 2. Existing LLM-based Methods for Vulnerability Detection
LLMs LLM Adaptation Techniques Data Deployment
Size
Data-
centric
LLM+
PA
LLM+
DL
Domain
Training
Causal
Learning
Default
Tuning
Prompt
Eng.
RAG
Input
Granularity
Data
Labeling
Interact
with Users
Integrated
to Workflow
Ziems et al. (DBLP:conf/infocom/ZiemsW21, ) ¡1B Function synthetic
Fu et al. (DBLP:conf/msr/FuT22, ) ¡1B Line heuristic
Hanif et al. (DBLP:conf/ijcnn/HanifM22, ) ¡1B Function
synthetic,
heuristic,
partially
manually labeled
Thapa et al. (DBLP:conf/acsac/ThapaJACPN22, ) ¡10B Function heuristic
Fu et al. (fu2023chatgpt, ) ¿10B Function heuristic
Wen et al. (DBLP:conf/kbse/WenWGWLG23, ) ¡1B Function
heuristic,
partially
manually labeled
Purba et al. (DBLP:conf/issre/PurbaGRC23, ) ¿1B Function heuristic
Liu et al. (DBLP:conf/dsc/LiuLGG23, ) ¿10B Function
partially
manually labeled
Peng et al. (DBLP:conf/scam/PengCZTLH23, ) ¡1B Line heuristic
Kuang et al. (DBLP:conf/scam/KuangYZTY23, ) ¡1B Function
heuristic,
partially
manually labeled
Zhang et al. (DBLP:journals/tse/ZhangLHXL23, ) ¡1B Function
heuristic,
partially
manually labeled
Tang et al. (DBLP:journals/jss/TangTBZF23, ) ¡1B Function
partially
manually labeled
Ni et al. (ni2023distinguishing, ) ¡1B Function
heuristic,
crowd-sourced
Zhou et al. (DBLP:journals/corr/abs-2401-15468, ) ¿10B Function heuristic
Fu et al. (DBLP:journals/ese/FuTLKNPG24, ) ¡1B Function heuristic
Sejfia et al. (DBLP:conf/icse/SejfiaDSM24, ) ¡1B Function
heuristic,
partially
manually labeled
Wang et al. (wang2023combining, ) ¡1B Function
synthetic,
heuristic,
manually labeled
Liu et al. (liu2024pre, ) ¡1B Function
heuristic,
partially
manually labeled
Rahman et al. (mahbubur2023towards, ) ¡1B Function
heuristic,
partially
manually labeled
Ding et al. (DBLP:journals/corr/abs-2403-18624, )
¡1B,
1-10B,
¿10B
Function
heuristic,
partially
manually labeled
Steenhoek et al. (DBLP:journals/corr/abs-2311-04109, ) ¡1B
Function,
Line
synthetic,
heuristic,
partially
manually labeled
Khare et al. (DBLP:journals/corr/abs-2311-16169, ) ¿1B Function
heuristic,
partially
manually labeled
Zhang et al. (DBLP:journals/corr/abs-2308-12697, ) ¿10B Function
synthetic,
heuristic
Risse et al. (DBLP:conf/uss/RisseB24, ) ¡1B Function
partially manually
labeled,
manually labeled
Steenhoek et al. (DBLP:conf/icse/SteenhoekRJL23, ) ¡1B
Function,
Line
heuristic
, partially
manually labeled
Yang et al. (DBLP:conf/icse/YangWLW23, ) ¡1B Function
heuristic,
partially
manually labeled
Chen et al. (DBLP:conf/raid/0001DACW23, ) ¡1B Function heuristic
Li et al. (DBLP:conf/icse/0027WZLZX0024, ) ¡1B Function heuristic
Croft et al. (DBLP:conf/icse/CroftBK23, ) ¡1B Function
heuristic,
manually labeled
Le et al. (DBLP:conf/msr/LeDB24, ) ¡1B
Function,
Line
heuristic,
partially
manually labeled
Zhou et al. (zhou2024comparison, )
¡10B
Repo
heuristic,
manually labeled
Tran et al. (tran2024detectvul, ) ¡1B Line heuristic
Jiang et al (DBLP:conf/internetware/JiangSGWW0024, ) ¡1B Function
heuristic,
partially
manually labeled
Shestov et al. (DBLP:journals/corr/abs-2401-17010, ) ¿10B Function
heuristic,
manually labeled
Ni et al (ni2024learning, ) ¿10B Function
heuristic
Weng et al. (DBLP:conf/internetware/WengQLLC24, ) ¡1B Line
heuristic
Du et al. (DBLP:journals/corr/abs-2406-11147, ) ¿10B Function
heuristic
Wen et al. (DBLP:journals/corr/abs-2404-15596, )
¡1B,
1-10B,
¿10B
Function
heuristic
Xin Yin (DBLP:journals/corr/abs-2404-03994, )
¿10B
Function
heuristic
Yang et al. (yang2024security, )
¡1B,
¿10B
Function
heuristic
In the following subsection, we will introduce detailed categories of LLM adaptation techniques for each of the following: 1) fine-tuning, 2) prompt engineering, and 3) retrieval augmentation (RAG). Table 2 summarizes the existing LLM-based methods for vulnerability detection.
5.1. Fine-tuning
Fine-tuning is a commonly used technique for adapting LLMs to vulnerability detection. In this process, labeled code samples, indicating whether they are vulnerable or not, are provided as training data. The model is then fine-tuned through supervised learning, where its parameters are adjusted based on these labeled examples. Several vulnerability detection studies (DBLP:conf/msr/FuT22, ; DBLP:conf/acsac/ThapaJACPN22, ; DBLP:journals/ese/FuTLKNPG24, ; DBLP:conf/icse/SejfiaDSM24, ; DBLP:conf/uss/RisseB24, ; DBLP:conf/icse/SteenhoekRJL23, ; DBLP:conf/raid/0001DACW23, ; DBLP:conf/icse/0027WZLZX0024, ; DBLP:conf/icse/CroftBK23, ; DBLP:conf/msr/LeDB24, ; zhou2024comparison, ; DBLP:journals/corr/abs-2401-17010, ) have directly applied this default fine-tuning approach without introducing additional designs. In this subsection, we focus on studies that propose advanced fine-tuning techniques that go beyond the default approach for improved effectiveness.
@@ 5.1. Fine-tuning
Data-centric Innovations. Data-centric innovations focus on optimizing the vulnerability detection data used for fine-tuning LLMs. Prior studies (DBLP:conf/icse/YangWLW23, ; DBLP:conf/icse/CroftBK23, ; DBLP:conf/issta/NieLWWLW23, ; DBLP:conf/raid/0001DACW23, ; DBLP:conf/scam/KuangYZTY23, ) have found that existing vulnerability detection data can suffer from imbalanced label distribution, noisy or incorrect labels, and scarcity of labeled data. Researchers (DBLP:conf/icse/YangWLW23, ; DBLP:conf/kbse/WenWGWLG23, ; DBLP:conf/scam/KuangYZTY23, ; DBLP:journals/corr/abs-2403-18624, ) have explored how to address the data issues.
@@
  • Imbalanced Learning: To address label imbalance, i.e., having more non-vulnerable code samples than vulnerable ones in the dataset, Yang et al. (DBLP:conf/icse/YangWLW23, ) applied various data sampling techniques. In addition, Ding et al. (DBLP:journals/corr/abs-2403-18624, ) up-weighted the loss value for the rare class (i.e., the vulnerable samples) to pay comparable attention to both vulnerable and clean classes. They found that random oversampling on raw code data enhances the ability of the LLM-based vulnerability detection approach to learn real vulnerable patterns.
@@
  • Positive and Unlabeled Learning: To address label quality issues such as noisy or incorrect labels, Wen et al. (DBLP:conf/kbse/WenWGWLG23, ) proposed PILOT, which learns solely from positive (vulnerable) and unlabeled data for vulnerability detection. Specifically, PILOT generates pseudo-labels for selected unlabeled data and mitigates the data noise by using a mixed-supervision loss.
@@
  • Counterfactual Training: To enhance the diversity of labeled data, Kuang et al. (DBLP:conf/scam/KuangYZTY23, ) proposed perturbing user-defined identifiers in the source code while preserving the syntactic and semantic structure. This approach generates diverse counterfactual training data, which refers to hypothetical data (e.g., data after perturbing identifiers) differing from actual data (i.e., data without perturbation), useful for analyzing the effect of certain factors. Incorporating these counterfactual data enriches the training data for LLMs.
@@
Combination of LLM with Program Analysis. Many LLMs undergo pre-training on extensive datasets through unsupervised objectives like masked language modeling (devlin2018bert, ) or next token prediction (radford2019language, ). For those LLMs, they may prioritize capturing sequential features and could potentially overlook certain structural aspects crucial for understanding code. To address this limitation, several studies have proposed integrating program analysis techniques with LLMs. The idea involves utilizing program analysis to extract structural features/relations within code, which are then incorporated into LLMs to enhance their understanding. Specifically, Liu et al. (liu2024pre, ) leveraged Joern (Joern, ) to build the AST and PDG of the function, leveraging this data to pre-train their LLM to predict statement-level control dependencies and token-level data dependencies within the function. Peng et al. (DBLP:conf/scam/PengCZTLH23, ) utilized program slicing to extract control and data dependency information, aiding LLMs in vulnerability detection. Wang et al. (wang2023combining, ) proposed to learn program presentations by feeding static source code information and dynamic program execution traces with LLMs. Additionally, Zhang et al. (DBLP:journals/tse/ZhangLHXL23, ) proposed decomposing syntax-based Control Flow Graphs (CFGs) into multiple execution paths and feeding these paths to LLMs for vulnerability detection. Tran et al. (tran2024detectvul, ) utilize a program analysis tool to normalize all user-defined names, such as variables and functions, into symbolic names (e.g., VAR_1, VAR_2, and FUNC_1) to enhance the model’s robustness against adversarial attacks. Weng et al. (DBLP:conf/internetware/WengQLLC24, ) incorporated inter-statement data and control dependency information into an LLM by masking irrelevant attention scores based on the program dependency graph, which resulted in improved code representation.
@@
Combination of LLM with Other Deep Learning Modules. LLMs have their own inherent limitations. Firstly, most LLMs are based on the Transformer architecture, which primarily models sequential relations and features. Secondly, some LLMs (e.g., CodeBERT) impose restrictions on the length of input code snippets. For instance, the most frequently used LLM in vulnerability detection, CodeBERT, can only process 512 tokens. To address these limitations, researchers have attempted to combine LLMs with other DL modules:
  • LLM+GNN: To leverage the structural features of code more effectively, Tang et al. (DBLP:journals/jss/TangTBZF23, ) introduced CSGVD, which employs graph neural networks (GNN) to extract the graph features of code and combine them with the features extracted by CodeBERT. Jiang et al. (DBLP:conf/internetware/JiangSGWW0024, ) parsed data flow graphs (DFGs) from the input code and used the data types of each variable as the node characteristics of the DFG. They then embedded the DFG information into the graph representation using a graph neural network (GNN) and integrated this graph representation with a large language model. In addition, Yang et al. (yang2024security, ) combined an LLM with a GNN model. Specifically, they first used data flow analysis (DFA) embeddings to establish a GNN model. They then concatenated the learned embeddings from the GNN with the hidden states of the LLM. By concatenating the embeddings during each forward pass, they enabled simultaneous training of both the LLM and the GNN.
@@
  • LLM+Bi-LSTM: To address the length constraints of LLMs, Ziems et al. (DBLP:conf/infocom/ZiemsW21, ) first segmented the input code into multiple fixed-size segments. They then utilized BERT to encode each segment and incorporated a Bidirectional Long Short-Term Memory (Bi-LSTM) module to process the output of BERT on each segment. Finally, a softmax classifier was applied to the last hidden state of the Bi-LSTM to produce the final classification scores.
@@
Domain-specific Pre-training. Domain-specific pre-training involves pre-training an LLM on data specific to a particular domain, such as vulnerability data, before fine-tuning it for a specific task within that domain. This process enables the LLM to better understand the data relevant to the domain. Several studies in the field of vulnerability detection have adopted this technique (DBLP:conf/ijcnn/HanifM22, ; liu2024pre, ; ni2023distinguishing, ; wang2023combining, ; DBLP:journals/corr/abs-2403-18624, ; DBLP:journals/corr/abs-2311-04109, ). These studies typically use one of three pre-training objectives:
@@
  • Masked Language Modeling: This pre-training objective involves training the LLM to predict masked tokens in a corrupted code snippet. Hanif and Mahmood (DBLP:conf/ijcnn/HanifM22, ) introduced VulBERTa, which pre-trained a RoBERTa (liu2019roberta, ) model on open-source C/C++ projects using the Masked Language Modeling objective. Since C/C++ is the programming language of the vulnerability detection data used to evaluate VulBERTa, this pre-training enhances VulBERTa’s ability to understand C/C++ code.
@@
  • Contrastive Learning: This pre-training objective is to minimize the distance between similar functions while maximizing the distance between dissimilar functions. Through contrastive learning, the LLM can learn to capture the distinctive features of code. In particular, Ni et al. (ni2023distinguishing, ) and Wang et al. (wang2023combining, ) utilized different hidden dropout masks to transform the same input function into positive samples (i.e., those similar to the input function) while regarding other distinct functions as dissimilar samples. Then they pre-trained an LLM following the contrastive learning objective. Additionally, Ding et al. (DBLP:journals/corr/abs-2403-18624, ) performed Class-aware Contrastive Learning (CA-CLR) to maximize the representation similarity between each sample and the perturbed version of itself and minimize the representation similarity between two randomly chosen samples.
@@
  • Predicting Program Dependencies: This pre-training objective aims to enhance the LLM by guiding the model to learn the knowledge necessary for analyzing dependencies in programs. Specifically, Liu et al. (liu2024pre, ) pre-trained an encoder-only LLM on code from open-source C/C++ projects to predict the statement-level control dependency and token-level data dependency.
  • Annotating Vulnerable Statements: This pre-training objective assumes that improving alignment to potentially vulnerable statements (PVS) would improve the performance of the model. For example, Steenhoek et al. (DBLP:journals/corr/abs-2311-04109, ) developed two annotation approaches (i.e., Mark and Prepend) to give guidance or extra information to the model. Mark inserted special “marker” tokens before and after each token inside the PVS, while Prepend added the tokens inside PVS at the beginning of the code. This can be considered as inserting domain knowledge into the model’s input.
@@
Causal Learning. Although promising, LLMs have been found to lack robustness under perturbation or when encountering out-of-distribution (OOD) data (mahbubur2023towards, ). Rahman et al. (mahbubur2023towards, ) suggested that this weak robustness may be attributed to LLMs learning non-robust features, such as variable names, that have spurious correlations with labels. To address this issue, Rahman et al. proposed CausalVul, which first designs perturbations to identify spurious features and then applies causal learning algorithms, specifically do-calculus, on top of LLMs to promote causal-based prediction, enhancing the robustness of LLMs in vulnerability detection.
5.2. Prompt Engineering.
Zero-shot Prompting. Researchers have attempted to devise effective prompts to guide LLMs in conducting vulnerability detection in the zero-shot prompting manner. Their prompt engineering designs generally consist of one or more components outlined below:
  • Task Description. This part of the prompt aims to equip LLMs with valuable task-specific information related to vulnerability detection, clarifying the objectives the model should achieve (e.g., identifying vulnerabilities in the provided code). Different studies employ varying task descriptions, and there is no consensus on which approach is optimal. Specifically, Zhou et al. (DBLP:journals/corr/abs-2401-15468, ) used a task description: “If has any potential vulnerability, output: ‘this code is vulnerable’. Otherwise, output: ‘this code is non-vulnerable’. The code is [code]. Let’s start:”. Fu et al. (fu2023chatgpt, ) proposed a task description: “Predict Whether the C/C++ function below is vulnerable. Strictly return 1 for a vulnerable function and 0 for a non-vulnerable function without further explanation.” Purba et al. (DBLP:conf/issre/PurbaGRC23, ) devised a task description: “[code] Is this code vulnerable? Answer in only Yes or No”. “[code]” refers to the input code in these task descriptions. Xin Yin (DBLP:journals/corr/abs-2404-03994, ) used the task description: “I will provide you a C code snippet and want you to tell whether it has a vulnerability. You need to output “yes” or “no” first (output no if uncertain), and then explain.” Zhou et al. (zhou2024comparison, ) used “If the following code snippet has any vulnerabilities, output Yes; otherwise, output No”.
  • Role Description. This aims to help LLMs shift their operational mode from a general language model to that of a vulnerability detector. Specifically, Zhou et al. (DBLP:journals/corr/abs-2401-15468, ) defined a role description: “You are an experienced developer who knows the security vulnerability very well”. Fu et al. (fu2023chatgpt, ) used a role description: “I want you to act as a vulnerability detection system”. Khare et al. (DBLP:journals/corr/abs-2311-16169, ) utilized a role description: You are a security researcher”. Xin Yin (DBLP:journals/corr/abs-2404-03994, ) used the role description as: “I want you to act as a vulnerability scanner”.
  • Vulnerability-Related Auxiliary Information. Some studies have integrated vulnerability-related auxiliary information into their prompt designs to enhance LLM performance. Specifically, Zhou et al.(DBLP:journals/corr/abs-2401-15468, ) suggested including vulnerable code examples representing the top 25 most dangerous CWEs of 2022 as part of the prompts to help LLMs better understand the characteristics of serious vulnerabilities. Additionally, Khare et al.(DBLP:journals/corr/abs-2311-16169, ) specified the types of vulnerabilities when querying LLMs about potential vulnerabilities in the provided code.
  • Program-Analysis Auxiliary Information. Some studies utilize program-analysis tools to extract code dependencies from the input code, aiming to enhance LLMs’ understanding of the input code. Specifically, Khare et al. (DBLP:journals/corr/abs-2311-16169, ) designed a dataflow analysis-based prompt that guides the model to simulate a source-sink-sanitizer-based dataflow analysis on the target code snippet before predicting if it is vulnerable. Zhang et al. (DBLP:journals/corr/abs-2308-12697, ) designed prompts that incorporate auxiliary information on data flow or API calls.
  • Chain-of-thought (CoT) Prompting. It is proposed by Kojima et al. (DBLP:conf/nips/KojimaGRMI22, ) for improved reasoning, which involves adding “Let’s think step by step” to the original prompt. Specifically, Zhou et al. (zhou2024comparison, ) and Ni et al. (ni2024learning, ) added the “Let’s think step by step” into the prompt.
Few-shot Prompting. In this approach, a few examples of input and ground truth label pairs are provided to the large language models (LLMs) as additional guidance. Specifically, Ni et al. (ni2024learning, ) utilize between 1 and 6 few-shot learning examples to fill the fixed context window (i.e., 4,096 tokens). Zhou et al. (zhou2024comparison, ) use two few-shot examples.
5.3. Retrieval Augmentation (RAG).